OpenStudy (anonymous):

Solve the following equation for x round the answer to four decimal places. 28^x = 10^-3x

OpenStudy (anonymous):

\[\LARGE 28^x=10^{-3x}\] \[\LARGE 28^x=\frac{1}{10^{3x}}\] \[\LARGE 28^x\times 10^{3x }=1\] \[\LARGE 28^x\times 1000^x=1\] \[\LARGE 28000^x=1\] and now I think you know the answer ...

OpenStudy (anonymous):

zero

OpenStudy (anonymous):

that's right... @dark_knight

OpenStudy (anonymous):

Why is the answer 0?